ABC Company's stock has a beta of 1.95, the risk-free rate is 2.25%, and the market risk premium is 6.75%. What is ABC's required rate of return using CAPM? Enter your answer rounded to two decimal places. Do not enter % in the answer box. For example, if your answer is 0.12345 or 12.345% then enter as 12.35 in the answer box. Ripken Iron Works believes the following probability distribution exists for its stock. What is the standard deviation of return on the company's stock? Enter your answer rounded to two decimal places. Do not enter % in the answer box. For example, if your answer is 0.12345 or 12.345% then enter as 12.35 in the answer box.

State of the Economy Probability of State Occurring Stock's Expected Return Boom 0.25 35% Normal 0.45 13% Recession 0.30 -26% Joel Foster is the portfolio manager of the Go Anywhere Fund, a $3 million hedge fund that contains the following stocks. The required rate of return on the market is 8.75% and the risk-free rate is 2.50%. What rate of return should investors expect (and require) on this fund? Enter your answer rounded to two decimal places. Do not enter % in the answer box. For example, if your answer is 0.12345 or 12.345% then enter as 12.35 in the answer box.

Stock Amount Beta

A $1,075,000 1.20

B 675,000 1.50

C 750,000 2.55

D 500,000 1.10

$3,000,000

1. Hazel Morrison, a mutual fund manager, has a $60 million portfolio with a beta of 1.00. The risk-free rate is 3.25%, and the market risk premium is 6.00%. Hazel expects to receive an additional $40 million, which she plans to invest in additional stocks. After investing the additional funds, she wants the fund's required and expected return to be 14%. What must the average beta of the new stocks be to achieve the target required rate of return? Enter your answer rounded to two decimal places. For example, if your answer is 123.45% or 1.2345 then enter as 1.23 in the answer box.
Answer #1
1) ABC COMPANY:
Required return using CAPM = Risk free rate+Beta*Market risk premium = 2.25%+1.95*6.75% = 15.41%
2) RIPKEN IRON WORKS:
Economy Probability [p] Return [r] E[r] = [p*r] Deviations [d] = r-E[r] d^2 p*d^2
Boom 0.25 35 8.75 28.20 795.24 198.81
Normal 0.45 13 5.85 6.20 38.44 17.298
Recession 0.30 -26 -7.80 -32.80 1075.84 322.752
6.80 538.86
SD of return = 538.86^0.5 = 23.21%
3) GO ANYWHERE FUND:
The first step is to find the weighted average beta of the stocks in the fund.
Weighted average beta of the stocks = Beta of the fund
Stock Amount Weight Beta Wt avg Beta
A $ 10,75,000 35.83% 1.20 0.43
B $    6,75,000 22.50% 1.50 0.34
C $    7,50,000 25.00% 2.55 0.64
D $    5,00,000 16.67% 1.10 0.18
Total $ 30,00,000 100.00% 1.59
Required return using CAPM = Risk free rate+Beta*(Market return-risk free rate)
= 2.50%+1.59*(8.75%-2.50%) = 12.44%
HAZEL MORRISON:
The first step is to find the beta for which the required return will be 14%.
So, 14 = 3.25+beta*6
beta = (14-3.25)/6 = 1.79
The beta of the revised portfolio should be 1.79.
That is, 1.79 should be the weighted average of
the beta of the existing portfolio and the beta
of the new stocks added.
That is, 1.79 = 1*60/100+b1*40/100, where b1 is
the beta of the new stocks.
Solving for b1
b1 = (1.79-0.6)*100/40 = 2.98